OpenStudy (anonymous):

(1+ cos theta) (1-cos theta) = sin^2 theta

OpenStudy (anonymous):

expanding...\[1 - \cos \theta + \cos \theta -\cos^2 \theta \rightarrow 1 - \cos^2 \theta = \sin^2 \theta \]

OpenStudy (anonymous):

is cos^2 theta + sin \^2 theta the answer?